# SpeeDe3DGS: Speedy Deformable 3D Gaussian Splatting with Temporal Pruning and Motion Grouping

### v3

@@ 5.1 SpeeDe3DGS Results
We first evaluate the components of SpeeDe3DGS on the real-world NeRF-DS dataset [yan2023nerf] in Table 1. Applying TSP alone accelerates rendering by , reduces the number of Gaussians by , and shortens training time by on average, while maintaining the image quality of DeformableGS [yang2023deformable3dgs]. Adding TSS further improves PSNR and SSIM beyond the baseline, as temporal perturbations act as a regularizer that suppresses floaters and stabilizes motion. Qualitative comparisons in Figure 3 show that TSP and TSS together effectively remove floaters and improve temporal coherence without sacrificing sharpness.
@@
Additional experiments on the synthetic D-NeRF [pumarola2020dnerf] and real-world HyperNeRF [park2021hypernerf] datasets are provided in Appendix A.2. On HyperNeRF, TSP+TSS and GroupFlow individually achieve and rendering speedups, respectively. When used together, they reach faster rendering, Gaussian reduction, and shorter training time, demonstrating the synergy of our SpeeDe3DGS methods. Training times for NeRF-DS and D-NeRF are measured on RTXA5000 GPUs (24 GB). We use RTXA6000 GPUs (48 GB) to train HyperNeRF due to the baseline memory requirements; however, pruning enables efficient training on 24 GB of VRAM.
@@
| EffGS [katsumata2023efficient] | 21.84 | 0.672 | 0.725 | 0.347 | 177.21 | 3757.81 |
| STG-decoder [li2023spacetime] | 21.81 | 0.678 | 0.742 | 0.352 | 109.42 | 5980.64 |
| STG [li2023spacetime] | 19.51 | 0.583 | 0.643 | 0.475 | 181.70 | 5359.56 |
| RTGS [yang2023ICLR4dgs] | 21.61 | 0.663 | 0.720 | 0.350 | 143.37 | 7352.52 |
| 4DGS [Wu2024CVPR4dgs] | 23.55 | 0.708 | 0.765 | 0.277 | 62.99 (1.00) | 8628.89 (1.00) |
@@ 5.2 MonoDyGauBench Evaluation
We evaluate our approach on the 50 scenes in MonoDyGauBench [liang2025monocular], extending the high-fidelity yet computationally heavy DeformableGS [yang2023deformable3dgs] and 4DGS [Wu2024CVPR4dgs] baselines with our Pruning (TSP+TSS) and GroupFlow modules. Table 2 summarizes aggregate results across all scenes, while Appendix A.1 provides per-dataset breakdowns. Qualitative results are presented in Appendix A.6. Although MonoDyGauBench experiments are nominally conducted on an RTX 3090 GPU, several baseline scenes exceed its 24 GB memory limit – a trend also noted in Section 5.1. Our pruning integration markedly reduces memory usage, enabling training within 24 GB on an RTX A5000 GPU.
As shown in Table 2, pruning makes 4DGS on-par with the fastest non-neural baselines while maintaining high image quality across metrics. Adding GroupFlow on top of pruning achieves the highest rendering speed – over 100 FPS faster than the fastest baseline – but with some loss of image quality, suggesting that per-Gaussian inference in DeformableGS better preserves fidelity under grouped motion.
Integrating pruning into DeformableGS accelerates rendering by , reaching frame rates comparable to RTGS [lin2024rtgs] while maintaining high visual fidelity across all datasets. It surpasses all non-neural baselines and our 4DGS variants in overall image quality, reinforcing DeformableGS as a strong backbone for neural motion modeling. Adding GroupFlow on top of pruning further increases rendering speed to , making it nearly 100 FPS faster than all baseline methods and on-par with our 4DGS+Pruning+GroupFlow variant. Training time is also the shortest, reduced by . Despite these speedups, reconstruction quality remains high: SSIM and MS-SSIM exceed the original DeformableGS baseline, suggesting a mild regularizing effect from GroupFlow. As discussed in Appendix A.1, this effect is most pronounced in scenes with unstable pose estimation. Overall, our SpeeDe3DGS strategies achieve an effective balance between speed, fidelity, and efficiency for dynamic Gaussian Splatting.
@@ 6 Limitations
While SpeeDe3DGS achieves substantial acceleration with minimal fidelity loss, some trade-offs remain. Minor quality degradation can occur under extreme pruning ratios, though this can be mitigated by adjusting pruning hyperparameters to balance speed and fidelity. Similarly, GroupFlow preserves quality across most scenes by modeling locally rigid motion, but highly deformable or non-rigid regions may experience reduced fidelity when the number of motion groups is limited. We further analyze these trade-off through ablations of pruning percentages and number of groups in Appendix A.4 and A.3. Our methods also introduce mild regularizing effects that improve temporal stability but do not incorporate explicit motion priors or learned dynamics; however, they are fully complementary to such approaches and can serve as acceleration modules for prior-driven or motion-aware 3DGS frameworks. In practice, TSP and TSS can be applied to any dynamic 3DGS framework, while GroupFlow extends to those with an explicit motion model. For a broader analysis of dynamic Gaussian Splatting methods and their limitations, we refer readers to MonoDyGauBench [liang2025monocular].
@@ 7 Conclusion
We introduced Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS), a framework that bridges the efficiency–fidelity gap in dynamic 3D Gaussian Splatting by reducing redundant neural inference through pruning and motion distillation. Our Temporal Sensitivity Pruning (TSP) module removes redundant Gaussians via temporally aggregated sensitivity analysis, while Temporal Sensitivity Sampling (TSS) stabilizes this process by perturbing timestamps to suppress floaters and enhance temporal coherence. Complementing pruning, GroupFlow distills neural motion into shared rigid SE(3) transformations, enabling efficient groupwise deformation and faster convergence. Across the fifty dynamic scenes in MonoDyGauBench [liang2025monocular], pruning accelerates DeformableGS by in rendering speed with minimal fidelity loss and fewer primitives, and adding GroupFlow further improves performance to faster rendering and shorter training while preserving superior image quality. Together, these modules demonstrate that temporally aware pruning and groupwise motion distillation can deliver neural-field fidelity at real-time speeds.
@@ A.1 Per-Dataset Metrics for MonoDyGauBench
We analyze results on Monocular Dynamic Gaussian Splatting Benchmark (MonoDyGauBench) [liang2025monocular], which evaluates dynamic 3DGS variants across five datasets. For consistency, we directly use the baseline tables from the manuscript. Both 4DGS and DeformableGS – the neural motion-field baselines – are extended with our pruning (TSP+TSS) and GroupFlow modules. Performance differences with our SpeeDe3DGS framework arise because MonoDyGauBench standardizes all methods within a unified training wrapper, integrating every motion representation and our components under common hyperparameters. Although MonoDyGauBench experiments are nominally conducted on an RTX 3090 GPU, baseline scenes in the HyperNeRF, Nerfies, and iPhone datasets exceed its 24 GB memory limit. Our pruning integration markedly reduces memory usage, enabling training within 24 GB on an RTX A5000 GPU. Qualitative results are presented in Appendix A.6.
On D-NeRF [pumarola2020dnerf], Table 3 shows that pruning and GroupFlow preserve the high neural fidelity of both baselines, surpassing all non-neural methods by a wide margin. Interestingly, adding GroupFlow slightly decreases rendering and training speed for DeformableGS – a behavior unique to D-NeRF within MonoDyGauBench and not observed in our SpeeDe3DGS framework results in Appendix A.2.
On HyperNeRF [park2021hypernerf], Table 4 shows that adding pruning to DeformableGS improves PSNR and SSIM beyond the baseline with an rendering speedup, while adding GroupFlow further enhances SSIM and MS-SSIM, culminating in a speedup. Efficiency improvements with our SpeeDe3DGS framework on HyperNeRF are reported in Appendix A.2.
On NeRF-DS [yan2023nerf], Table 5 shows that adding TSP and TSS yields comparable or improved image quality over the DeformableGS baseline, while adding GroupFlow provides a rendering acceleration with minimal fidelity loss. These trends are consistent with our findings in Section 5.1. The 4DGS baseline performs comparatively poorly, consistent with MonoDyGauBench’s broader analysis of motion–representation trade-offs.
On Nerfies [park2021nerfies], Table 6 shows that pruning improves all quality metrics for DeformableGS, demonstrating the stabilizing effects of TSP+TSS. Adding GroupFlow further increases efficiency, achieving a rendering speedup and the highest FPS in the table, while maintaining higher image quality than all baselines.
On iPhone [gao2022dynamic], Table 7 shows that adding GroupFlow significantly improves PSNR, SSIM, and MS-SSIM while achieving a rendering speedup. By enforcing locally consistent motion through shared rigid SE(3) transformations, GroupFlow acts as a strong structural regularizer that enhances temporal coherence and mitigates drift under noisy or unstable camera trajectories.
@@
| EffGS [katsumata2023efficient] | 30.52 | 0.96 | 0.98 | 0.04 | 289.47 | 1042.04 |
| STG-decoder [li2023spacetime] | 25.89 | 0.91 | 0.90 | 0.17 | 160.32 | 3462.29 |
| STG [li2023spacetime] | 17.09 | 0.88 | 0.66 | 0.29 | 208.98 | 4889.50 |
| RTGS [yang2023ICLR4dgs] | 28.78 | 0.96 | 0.96 | 0.05 | 192.37 | 1519.60 |
| 4DGS [Wu2024CVPR4dgs] | 33.27 | 0.98 | 0.99 | 0.02 | 134.13 (1.00) | 1781.46 (1.00) |
@@
| EffGS [katsumata2023efficient] | 22.24 | 0.70 | 0.79 | 0.37 | 138.79 | 4119.66 |
| STG-decoder [li2023spacetime] | 23.92 | 0.73 | 0.83 | 0.34 | 66.72 | 7423.41 |
| STG [li2023spacetime] | 22.92 | 0.70 | 0.78 | 0.41 | 183.21 | 5729.80 |
| RTGS [yang2023ICLR4dgs] | 22.99 | 0.71 | 0.79 | 0.35 | 104.64 | 11507.80 |
| 4DGS [Wu2024CVPR4dgs] | 25.70 | 0.79 | 0.89 | 0.23 | 37.46 (1.00) | 10170.86 (1.00) |
@@
| EffGS [katsumata2023efficient] | 21.28 | 0.78 | 0.77 | 0.25 | 307.70 | 1597.90 |
| STG-decoder [li2023spacetime] | 21.73 | 0.80 | 0.81 | 0.20 | 212.24 | 2131.48 |
| STG [li2023spacetime] | 20.13 | 0.69 | 0.72 | 0.39 | 302.89 | 2214.62 |
| RTGS [yang2023ICLR4dgs] | 19.88 | 0.75 | 0.73 | 0.30 | 259.83 | 3116.21 |
| 4DGS [Wu2024CVPR4dgs] | 20.08 | 0.73 | 0.71 | 0.25 | 100.22 (1.00) | 4075.43 (1.00) |
@@
| EffGS [katsumata2023efficient] | 20.13 | 0.43 | 0.61 | 0.65 | 159.17 | 3249.33 |
| STG-decoder [li2023spacetime] | 20.93 | 0.47 | 0.67 | 0.58 | 90.46 | 6050.42 |
| STG [li2023spacetime] | 20.55 | 0.46 | 0.66 | 0.62 | 142.30 | 5264.58 |
| RTGS [yang2023ICLR4dgs] | 20.06 | 0.42 | 0.62 | 0.61 | 153.38 | 9059.54 |
| 4DGS [Wu2024CVPR4dgs] | 21.84 | 0.50 | 0.72 | 0.47 | 35.70 (1.00) | 9797.50 (1.00) |
@@
| EffGS [katsumata2023efficient] | 16.82 | 0.47 | 0.50 | 0.47 | 99.64 | 6275.33 |
| STG-decoder [li2023spacetime] | 16.85 | 0.47 | 0.52 | 0.49 | 86.18 | 7694.83 |
| STG [li2023spacetime] | 15.60 | 0.40 | 0.45 | 0.57 | 114.95 | 6629.64 |
| RTGS [yang2023ICLR4dgs] | 15.36 | 0.39 | 0.44 | 0.52 | 101.33 | 8484.05 |
| 4DGS [Wu2024CVPR4dgs] | 15.13 | 0.42 | 0.43 | 0.52 | 42.53 (1.00) | 14205.48 (1.00) |
@@ A.2 Additional SpeeDe3DGS Results
We extend our analysis to the synthetic D-NeRF [pumarola2020dnerf] and real-world HyperNeRF [park2021hypernerf] datasets using the DeformableGS [yang2023deformable3dgs] codebase. Rendering FPS is measured on RTX 3090 GPUs with 24GB of VRAM for consistency with MonoDyGauBench[liang2025monocular].
On D-NeRF [pumarola2020dnerf], both pruning and GroupFlow individually provide around a increase in rendering speed and a reduction in training time, measured on RTXA5000 GPUs (24GB). Pruning slightly reduces reconstruction quality more than GroupFlow, likely because D-NeRF’s synthetic scenes are compact, noise-free, and already sparsely parameterized, leaving limited redundancy to remove. Nevertheless, pruning achieves a reduction in Gaussian count. Since these scenes contain no pose noise or floaters, TSP and TSP+TSS perform nearly identically, showing that TSS adds stability without affecting clean synthetic settings. GroupFlow performs particularly well due to the simple, rigid motions of D-NeRF scenes. When combined, pruning and GroupFlow achieve a rendering speedup and a reduction in training time.
For HyperNeRF [park2021hypernerf], we report results on the same eight real-world scenes used in DeformableGS – espresso, americano, cookie, chicken, torchocolate, lemon, hand, and printer. Following DeformableGS, image-quality metrics are not reported because most scenes lack evaluation splits; we provide full results on MonoDyGauBench in Table 4. Pruning alone yields a rendering speedup, an reduction in Gaussians, and a reduction in training time. GroupFlow independently achieves a rendering speedup and a shorter training time. Together, they reach faster rendering and shorter training, demonstrating strong synergy between our SpeeDe3DGS components. Baseline HyperNeRF models require RTX A6000 GPUs with 48GB of VRAM for training, whereas our pruned configurations fit comfortably within 24GB on RTX A5000s. All inference metrics, including FPS, are collected on RTX3090 GPUs (24GB), matching the MonoDyGauBench setup and underscoring that baseline models exceed 24 GB memory during training despite being evaluated on 3090s for inference.

### v4

@@ 5.1 SpeeDe3DGS Results
We first evaluate the components of SpeeDe3DGS on the real-world NeRF-DS dataset [52] in Table 1. Applying TSP alone accelerates rendering by , reduces the number of Gaussians by , and shortens training time by on average, while maintaining the image quality of DeformableGS [54]. Adding TSS further improves PSNR and SSIM beyond the baseline, as temporal perturbations act as a regularizer that suppresses floaters and stabilizes motion. Qualitative comparisons in Figure 3 show that TSP and TSS together effectively remove floaters and improve temporal coherence without sacrificing sharpness.
@@
Results on the synthetic D-NeRF [39] and real-world HyperNeRF [37] datasets are provided in Appendix A.2. On HyperNeRF, TSP + TSS and GroupFlow individually achieve and rendering speedups, respectively. When used together, they reach faster rendering, Gaussian reduction, and shorter training time, demonstrating the synergy of our methods. Training times for NeRF-DS and D-NeRF are measured on RTX A5000 GPUs (24 GB). We use RTX A6000 GPUs (48 GB) to train HyperNeRF due to the baseline memory requirements; however, pruning enables efficient training on 24 GB of VRAM. Per-scene metrics for all datasets are provided in Appendix A.6.
@@
| EffGS [19] | 21.84 | 0.672 | 0.725 | 0.347 | 177.21 | 3757.81 |
| STG-decoder [24] | 21.81 | 0.678 | 0.742 | 0.352 | 109.42 | 5980.64 |
| STG [24] | 19.51 | 0.583 | 0.643 | 0.475 | 181.70 | 5359.56 |
| RTGS [53] | 21.61 | 0.663 | 0.720 | 0.350 | 143.37 | 7352.52 |
| 4DGS [49] | 23.55 | 0.708 | 0.765 | 0.277 | 62.99 (1.00) | 8628.89 (1.00) |
@@ 5.2 MonoDyGauBench Evaluation
We evaluate our approach on the 50 scenes in MonoDyGauBench [25], extending the high-fidelity yet computationally heavy DeformableGS [54] and 4DGS [49] baselines with our Pruning (TSP + TSS) and GroupFlow modules. Table 2 summarizes aggregate results across all scenes, while Appendix A.1 provides per-dataset breakdowns. Qualitative results are presented in Appendix A.7. Although MonoDyGauBench experiments are nominally conducted on an RTX 3090 GPU, several baseline scenes exceed its 24 GB memory limit – a trend also noted in Section 5.1. Our pruning integration markedly reduces memory usage, enabling training within 24 GB on an RTX A5000 GPU.
As shown in Table 2, pruning makes 4DGS on par with the fastest non-neural baselines while maintaining high image quality across metrics. Adding GroupFlow on top of pruning achieves the highest rendering speed – over 100 FPS faster than the fastest baseline – but with some loss of image quality, suggesting that per-Gaussian inference in DeformableGS better preserves fidelity under grouped motion.
Integrating pruning into DeformableGS accelerates rendering by , reaching frame rates comparable to RTGS [26] while maintaining high visual fidelity across all datasets. It surpasses all non-neural baselines and our 4DGS variants in overall image quality, reinforcing DeformableGS as a strong backbone for neural motion modeling. Adding GroupFlow on top of pruning further increases rendering speed to , making it nearly 100 FPS faster than all baseline methods and on-par with our 4DGS+Pruning+GroupFlow variant. Training time is also the shortest, reduced by . Despite these speedups, reconstruction quality remains high: SSIM and MS-SSIM exceed the original DeformableGS baseline, suggesting a mild regularizing effect from GroupFlow. As discussed in Appendix A.1, this effect is most pronounced in scenes with unstable pose estimation. Overall, our SpeeDe3DGS strategies achieve an effective balance between speed, fidelity, and efficiency for dynamic Gaussian Splatting.
@@ 6 Limitations
While SpeeDe3DGS achieves substantial acceleration with minimal fidelity loss, some trade-offs remain. Minor quality degradation can occur under extreme pruning ratios, though this can be mitigated by adjusting pruning hyperparameters to balance speed and fidelity. Similarly, GroupFlow preserves quality in most scenes through locally rigid motion, but highly deformable regions may lose fidelity when the number of motion groups is limited. We analyze these trade-offs through ablations of pruning percentages and number of groups in Appendices A.5 and A.3. While our methods introduce mild regularizing effects that improve temporal stability, they do not incorporate explicit motion priors or learned dynamics. They are therefore complementary to prior-driven or motion-aware 3DGS frameworks and can accelerate such methods. In practice, TSP and TSS can be applied to any dynamic 3DGS framework, while GroupFlow extends to those with an explicit motion model. Appendix A.8 further discusses baseline and motion model limitations. For a broader analysis of dynamic Gaussian Splatting methods and their limitations, we refer readers to MonoDyGauBench [25].
@@ 7 Conclusion
We introduced Speedy Deformable 3D Gaussian Splatting (SpeeDe3DGS), a framework that bridges the efficiency–fidelity gap in dynamic 3D Gaussian Splatting by reducing redundant neural inference through pruning and motion distillation. Our Temporal Sensitivity Pruning (TSP) module removes redundant Gaussians via temporally aggregated sensitivity analysis, while Temporal Sensitivity Sampling (TSS) stabilizes this process by perturbing timestamps to suppress floaters and enhance temporal coherence. Complementing pruning, GroupFlow distills neural motion into shared rigid SE(3) transformations, enabling efficient groupwise deformation and faster convergence. Across the 50 dynamic scenes in MonoDyGauBench [25], pruning accelerates DeformableGS by in rendering speed with minimal fidelity loss and fewer primitives, and adding GroupFlow further improves performance to faster rendering and shorter training while preserving superior image quality. Together, these modules demonstrate that temporally aware pruning and groupwise motion distillation can deliver neural-field fidelity at real-time speeds.
@@ A.1 Per-Dataset Metrics for MonoDyGauBench
We analyze results on Monocular Dynamic Gaussian Splatting Benchmark (MonoDyGauBench) [25], which evaluates dynamic 3DGS variants across five datasets. For consistency, we directly use the baseline tables from the manuscript. Both 4DGS and DeformableGS – the neural motion-field baselines – are extended with our pruning (TSP+TSS) and GroupFlow modules. Performance differences with our SpeeDe3DGS framework arise because MonoDyGauBench standardizes all methods within a unified training wrapper, integrating every motion representation and our components under common hyperparameters. Although MonoDyGauBench experiments are nominally conducted on an RTX 3090 GPU, baseline scenes in the HyperNeRF, Nerfies, and iPhone datasets exceed its 24 GB memory limit. Our pruning integration markedly reduces memory usage, enabling training within 24 GB on an RTX A5000 GPU. Qualitative results are presented in Appendix A.7.
On D-NeRF [39], Table 3 shows that pruning and GroupFlow preserve the high neural fidelity of both baselines, surpassing all non-neural methods by a wide margin. Interestingly, adding GroupFlow slightly decreases rendering and training speed for DeformableGS – a behavior unique to D-NeRF within MonoDyGauBench and not observed in our SpeeDe3DGS framework results in Appendix A.2.
On HyperNeRF [37], Table 4 shows that adding pruning to DeformableGS improves PSNR and SSIM beyond the baseline with an rendering speedup, while adding GroupFlow further enhances SSIM and MS-SSIM, culminating in a speedup. Efficiency improvements with our SpeeDe3DGS framework on HyperNeRF are reported in Appendix A.2.
On NeRF-DS [52], Table 5 shows that adding TSP and TSS yields comparable or improved image quality over the DeformableGS baseline, while adding GroupFlow provides a rendering acceleration with minimal fidelity loss. These trends are consistent with our findings in Section 5.1. The 4DGS baseline performs comparatively poorly, consistent with MonoDyGauBench’s broader analysis of motion–representation trade-offs.
On Nerfies [36], Table 6 shows that pruning improves all quality metrics for DeformableGS, demonstrating the stabilizing effects of TSP+TSS. Adding GroupFlow further increases efficiency, achieving a rendering speedup and the highest FPS in the table, while maintaining higher image quality than all baselines.
On iPhone [10], Table 7 shows that adding GroupFlow significantly improves PSNR, SSIM, and MS-SSIM while achieving a rendering speedup. By enforcing locally consistent motion through shared rigid SE(3) transformations, GroupFlow acts as a strong structural regularizer that enhances temporal coherence and mitigates drift under noisy or unstable camera trajectories.
@@
| EffGS [19] | 30.52 | 0.96 | 0.98 | 0.04 | 289.47 | 1042.04 |
| STG-decoder [24] | 25.89 | 0.91 | 0.90 | 0.17 | 160.32 | 3462.29 |
| STG [24] | 17.09 | 0.88 | 0.66 | 0.29 | 208.98 | 4889.50 |
| RTGS [53] | 28.78 | 0.96 | 0.96 | 0.05 | 192.37 | 1519.60 |
| 4DGS [49] | 33.27 | 0.98 | 0.99 | 0.02 | 134.13 (1.00) | 1781.46 (1.00) |
@@
| EffGS [19] | 22.24 | 0.70 | 0.79 | 0.37 | 138.79 | 4119.66 |
| STG-decoder [24] | 23.92 | 0.73 | 0.83 | 0.34 | 66.72 | 7423.41 |
| STG [24] | 22.92 | 0.70 | 0.78 | 0.41 | 183.21 | 5729.80 |
| RTGS [53] | 22.99 | 0.71 | 0.79 | 0.35 | 104.64 | 11507.80 |
| 4DGS [49] | 25.70 | 0.79 | 0.89 | 0.23 | 37.46 (1.00) | 10170.86 (1.00) |
@@
| EffGS [19] | 21.28 | 0.78 | 0.77 | 0.25 | 307.70 | 1597.90 |
| STG-decoder [24] | 21.73 | 0.80 | 0.81 | 0.20 | 212.24 | 2131.48 |
| STG [24] | 20.13 | 0.69 | 0.72 | 0.39 | 302.89 | 2214.62 |
| RTGS [53] | 19.88 | 0.75 | 0.73 | 0.30 | 259.83 | 3116.21 |
| 4DGS [49] | 20.08 | 0.73 | 0.71 | 0.25 | 100.22 (1.00) | 4075.43 (1.00) |
@@
| EffGS [19] | 20.13 | 0.43 | 0.61 | 0.65 | 159.17 | 3249.33 |
| STG-decoder [24] | 20.93 | 0.47 | 0.67 | 0.58 | 90.46 | 6050.42 |
| STG [24] | 20.55 | 0.46 | 0.66 | 0.62 | 142.30 | 5264.58 |
| RTGS [53] | 20.06 | 0.42 | 0.62 | 0.61 | 153.38 | 9059.54 |
| 4DGS [49] | 21.84 | 0.50 | 0.72 | 0.47 | 35.70 (1.00) | 9797.50 (1.00) |
@@
| EffGS [19] | 16.82 | 0.47 | 0.50 | 0.47 | 99.64 | 6275.33 |
| STG-decoder [24] | 16.85 | 0.47 | 0.52 | 0.49 | 86.18 | 7694.83 |
| STG [24] | 15.60 | 0.40 | 0.45 | 0.57 | 114.95 | 6629.64 |
| RTGS [53] | 15.36 | 0.39 | 0.44 | 0.52 | 101.33 | 8484.05 |
| 4DGS [49] | 15.13 | 0.42 | 0.43 | 0.52 | 42.53 (1.00) | 14205.48 (1.00) |
@@ A.2 Additional SpeeDe3DGS Results
We extend our analysis to the synthetic D-NeRF [39] and real-world HyperNeRF [37] datasets using the DeformableGS [54] codebase. Rendering FPS is measured on an RTX 3090 GPU with 24 GB of VRAM for consistency with MonoDyGauBench [25].
On D-NeRF [39], both pruning and GroupFlow individually provide around a increase in rendering speed and a reduction in training time, measured on an RTX A5000 GPU (24 GB). Pruning slightly reduces reconstruction quality more than GroupFlow, likely because D-NeRF’s synthetic scenes are compact, noise-free, and already sparsely parameterized, leaving limited redundancy to remove. Nevertheless, pruning achieves a reduction in the number of Gaussians. Since these scenes contain no pose noise or floaters, TSP and TSP+TSS perform nearly identically, showing that TSS adds stability without affecting clean synthetic settings. GroupFlow performs particularly well due to the simple, rigid motions of D-NeRF scenes. When combined, pruning and GroupFlow achieve a rendering speedup and a reduction in training time.
For HyperNeRF [37], we report results on the same eight real-world scenes used in DeformableGS – espresso, americano, cookie, chicken, torchocolate, lemon, hand, and printer. Following DeformableGS, image-quality metrics are not reported because most scenes lack evaluation splits; we provide full results on MonoDyGauBench in Table 4. Pruning alone yields a rendering speedup, an reduction in Gaussians, and a reduction in training time. GroupFlow independently achieves a rendering speedup and a shorter training time. Together, they reach faster rendering and shorter training, demonstrating strong synergy between our SpeeDe3DGS components. Baseline HyperNeRF models require an RTX A6000 GPU with 48 GB of VRAM for training, whereas our pruned configurations fit comfortably within 24 GB on an RTX A5000. All inference metrics, including FPS, are collected on an RTX 3090 GPU (24 GB), matching the MonoDyGauBench setup and underscoring that baseline models exceed 24 GB memory during training despite being evaluated on an RTX 3090 for inference.